Calculate the mass, in grams, of iron(II) sulfate that must be added to a 500-mL volumetric flask in order to prepare 500 mL of a 0.196 M aqueous solution of the salt.


1
Expert's answer
2022-02-09T11:08:03-0500

"C= n\/v"

n = 0.196× 0.5= 0.098moles

Mass= 0.098× 152= 14.9g